We Pay Insurance Co. will pay you $1,450 each quarter for 27 years. You want to earn a minimum interest rate of .99 percent per quarter. What is the most you are willing to pay today for these payments?

$95,420.38

$91,124.72

$277,956.49

$92,723.40

$95,920.76
Answer #1

Installment per quarter for 27 years =$1450

Total Quarter =27*4 =108

Interest rate per quarter = 0.99%

Present value to pay for these payment would be = Installment* PVAF(.99%,108 period) = 1450*66.15 =$95920.76

$95,920.76 is correct answer.